OpenStudy (somegirl1999):

Find the values of x, y, and z. (1 point) x = 91°; y = 51°; z = 31° x = 89°; y = 91°; z = 0° x = 60°; y = 120°; z = 31° x = 91°; y = 89°; z = 31°

OpenStudy (somegirl1999):



OpenStudy (anonymous):

Angles in a triangle = 180. X = 180 - (38 + 51) Angles on a line = 180. Y = 180 - X So Z = 180 - (180-X + 60)

OpenStudy (somegirl1999):

thank you so much

OpenStudy (anonymous):

That's ok :)

OpenStudy (anonymous):

Is this not answered?

OpenStudy (somegirl1999):

Lol i meant to close it

OpenStudy (anonymous):

Ah fair enough :)

OpenStudy (somegirl1999):

Lol